### v2

# The Christoffel problem for the disk area measure

Abstract.
The mixed Christoffel problem asks for necessary and sufficient conditions for a Borel measure on the Euclidean unit sphere to be the mixed area measure of some convex bodies, all but one of them are fixed. We provide a solution in the case where the fixed reference bodies are (n1)(n-1)-dimensional disks with parallel axes and the measure has no mass at the poles of the sphere determined by this axis.
1. Introduction
For a convex body (convex, compact set) $K\subset\mathbb{R}^{n}$ with smooth boundary, the area measure of order $1$, denoted $S_{1}(K,\cdot)$, is the absolutely continuous measure on the unit sphere $\mathbb{S}^{n-1}$ with density given by the mean radius of curvature of $K$ as a function of the outer unit normal vector. Extending this definition, the definition of $S_{1}(K,\cdot)$ can be extended to general convex bodies via variations of the surface area measure $S_{n-1}(K,\cdot)$, that is,
(1.1) $S_{1}(K,\beta)=\frac{1}{(n-1)!}\left.\left(\frac{d}{dt}\right)^{\!n-2}\right|_{t=0^{+}}S_{n-1}(K+tB^{n},\beta),$
for every Borel set $\beta\subseteq\mathbb{S}^{n-1}$, where $B^{n}$ denotes the euclidean unit ball. Here, the surface area measure $S_{n-1}(K,\beta)$ evaluated at a Borel set $\beta$ is given by the Hausdorff measure $\mathcal{H}^{n-1}$ of all boundary points of $K$ with outer unit normal in $\beta$. We refer to [Schneider2014]*Sec. 4 for a more detailed exposition.
Dating back to the nineteenth century with the work of Christoffel [Christoffel1865], it was an important question in convex geometry, nowadays known as the Christoffel problem, to determine the class of Borel measures on the sphere that arise as $S_{1}(K,\cdot)$ for some convex body $K$. After Christoffel’s work [Christoffel1865], treating the three dimensional case, a complete solution was obtained at the end of the 1960s with the works of Berg [Berg1969] and Firey [Firey1967, Firey1968]. Special cases were treated for smooth bodies by Pogorelov [Pogorelov1953] and for polytopes by Schneider [Schneider1977b].
The key to the solution of the Christoffel problem was the observation that $S_{1}(K,\cdot)$ can be obtained from the support function $h_{K}$ of $K$ by the means of the differential operator
$\Box_{n}=\frac{1}{n-1}\Delta_{\mathbb{S}}+\mathrm{Id},$
where $\Delta_{\mathbb{S}}$ is the spherical laplacian on $\mathbb{S}^{n-1}$. Indeed, $S_{1}(K,\cdot)=\Box_{n}h_{K}$ in the distributional sense, and using spherical harmonics, Berg [Berg1969] constructed a family of functions $g_{n}$, $n\geq 2$, on (1,1)(-1,1) that serve as integral kernels of the inverse operator to $\Box_{n}$. Writing $\mathcal{K}(\mathbb{R}^{n})$ for the set of convex bodies in $\mathbb{R}^{n}$, Berg’s result [Berg1969]*Thm. 5.3 can be formulated as follows.
Theorem 1.1 ([Berg1969]).
Let $\mu$ be a centered, non-negative, finite Borel measure on $\mathbb{S}^{n-1}$. Then there exists a convex body $K\in\mathcal{K}(\mathbb{R}^{n})$ with $\mu=S_{1}(K,{{}\cdot{}})$ if and only if
(1.2) $h(u)=\int_{\mathbb{S}^{n-1}}g_{n}(\langle u,v\rangle)\,\mu(dv),\qquad u\in\mathbb{S}^{n-1},$
is a support function. In that case $h=h_{K-s(K)}$.
Here and in the following, a measure $\mu$ on $\mathbb{S}^{n-1}$ is called centered if $\int_{\mathbb{S}^{n-1}}ud\mu(u)=0$. Moreover, we denote by $\langle\cdot,\cdot\rangle$ the euclidean inner product and by
(1.3) $s(K)=\frac{1}{\kappa_{n}}\int_{\mathbb{S}^{n-1}}h_{K}(u)u\,du\in\mathbb{R}^{n}$
the Steiner point of a given convex body $K\in\mathcal{K}(\mathbb{R}^{n})$ (see, e.g., [Schneider2014]*p. 50).
Let us spell out the explicit formula for the function $g_{2}$ for later reference:
$g_{2}(t)=\sqrt{1-t^{2}}(\pi-\arccos t)+ct,$
with $c$ chosen in such a way to make the function $u\mapsto g_{2}(\langle e_{n},u\rangle)$ on $\mathbb{S}^{n-1}$ centered.
In this note, we consider an anisotropic version of $S_{1}(K,\cdot)$, obtained by replacing the unit ball $B^{n}$ in ˜1.1 by an arbitrary convex body $C\in\mathcal{K}(\mathbb{R}^{n})$
(1.4) $S_{1}(K,C;\beta)=\frac{1}{(n-1)!}\left.\left(\frac{d}{dt}\right)^{\!n-2}\right|_{t=0^{+}}S_{n-1}(K+tC,\beta),$
for every Borel set $\beta\subseteq\mathbb{S}^{n-1}$, which gives rise to the anisotropic Christoffel problem:
Problem.
Given $C\in\mathcal{K}(\mathbb{R}^{n})$, find necessary and sufficient conditions for a Borel measure $\mu$ on $\mathbb{S}^{n-1}$ such that $\mu=S_{1}(K,C;{}\cdot{})$ for some convex body $K\in\mathcal{K}(\mathbb{R}^{n})$.
This problem has only started to gain attention very recently (e.g., [Colesanti2025a]), and very little is known about it. Let us note that it was also considered in the broader range of an anisotropic Christoffel–Minkowski problem, that is, for intermediate area measures, in [Brauner2025a], where it was solved in the case where all bodies are bodies of revolution. Indeed, this note originated from a side observation during the project [Brauner2025a].
1.1. Main result
Our main result is a solution of the anisotropic Christoffel problem in the case where $C=\mathbb{D}$, that is, for the disk area measure $S_{1}(K,\mathbb{D};{{}\cdot{}})$, under an additional condition on the mass at the north and south pole $\pm e_{n}\in\mathbb{S}^{n-1}$. We hope that, in analogy to the proof strategy in [Brauner2025a], this can serve as an initial step to solve the general anisotropic Christoffel problem via transforming the area measures.
In order to state our theorem, let $\pi:\mathbb{S}^{n-1}\setminus\{\pm e_{n}\}\to\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$ denote the map that assigns to every $u\in\mathbb{S}^{n-1}\setminus\{\pm e_{n}\}$ the unique linear $2$-space $e_{n}\vee u$ containing both $u$ and the $n$-th coordinate vector $e_{n}$. Here, we denote by $\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$ the grassmanian of all $2$-dimensional subspaces of $\mathbb{R}^{n}$ that contain $e_{n}$.
Under the assumption that the pushforward $\pi_{\ast}\mu$ is absolutely continuous with continuous density, the measure $\mu$ disintegrates into a family $(\mu_{E})_{E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})}$ of measures on $S^{1}(E)$, such that
(1.5) $\int_{\mathbb{S}^{n-1}\setminus\{\pm e_{n}\}}f(u)\,\mu(du)=\int_{\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})}\int_{\mathbb{S}^{1}(E)\setminus\{\pm e_{n}\}}f(v)\,\mu_{E}(dv)\,dE.$
Note that, using the condition that $\mu(\{\pm e_{n}\})=0$, we tacitly identify $\mu$ with its restriction to $\mathbb{S}^{n-1}\setminus\{\pm e_{n}\}$. See Section˜2.2 for the details on the construction of this disintegration.
Theorem A.
Let $\mu$ be a non-negative, centered, finite Borel measure on $\mathbb{S}^{n-1}$ and suppose that $\mu(\{\pm e_{n}\})=0$. Then there exists a convex body $K\in\mathcal{K}(\mathbb{R}^{n})$ with $\mu=S_{1}(K,\mathbb{D},{}\cdot{})$if and only if
  1. (i)
    The measure $\pi_{\ast}\mu$ is absolutely continuous with a continuous density,
  2. (ii)
    $\int_{\mathbb{S}^{1}(E)}v\,\mu_{E}(dv)=0$ for a.e. $E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$, and
  3. (iii)
    there exists a support function $h$ of a convex body such that
    (1.6) $h(u)-\frac{1}{\pi}\int_{\mathbb{S}^{1}(e_{n}\vee u)}h(v)\langle u,v\rangle\,dv=\int_{\mathbb{S}^{1}(e_{n}\vee u)}\sqrt{1-\langle u,v\rangle^{2}}(\pi-\arccos\langle u,v\rangle)\,\mu_{e_{n}\vee u}(dv),$
    for a.e. $u\in\mathbb{S}^{n-1}$.
In this case, $K$ is unique up to a translation, and $h=h_{K-v}$ for some $v\in\mathbb{R}^{n}$.
Let us comment on the condition that the measure has no mass at the poles $\pm e_{n}$. This is due to the fact that the solution of the Christoffel problem for the disk is not unique as soon as there is a face in either direction $\pm e_{n}$. Indeed, for every $K\in\mathcal{K}(e_{n}^{\perp})$, $S_{1}(K,\mathbb{D},\cdot)=V^{e_{n}^{\perp}}(K,\mathbb{D}^{[n-2]})(\delta_{e_{n}}+\delta_{-e_{n}})$ only depends on the mean width of $K$.
On a technical level, mass at the poles $\pm e_{n}$ raises the problem of how to define a disintegration similar to ˜1.5. Indeed, the domains of the measures $\mu_{E}$ would overlap in the poles $\pm e_{n}$, whence a method is needed to distribute the mass $\mu(\{\pm e_{n}\})$ among the measures $\mu_{E}$. However, this corresponds to determine the (highly non-unique) shape of the faces of $K$ in direction $\pm e_{n}$. As soon as a disintegration satisfying ˜1.6 is constructed, the steps in the proof of Theorem˜A can be repeated to yield the solution.
2. Proof of the main result
Our proof is inspired by the the Kubota-type formula of Hug, Mussnig, and Ulivelli [Hug2024] in the instance where $i=1$: For a convex body $K\in\mathcal{K}(\mathbb{R}^{n})$ and a bounded Borel function $f:\mathbb{S}^{n-1}\to\mathbb{R}$,
(2.1) $\int_{\mathbb{S}^{n-1}}f(u)\,S_{1}(K,\mathbb{D},du)=\frac{\kappa_{n-1}}{2}\int_{\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})}\!\int_{\mathbb{S}^{1}(E)}\!f(u)\,S^{E}_{1}(K|E,du)\,dE.$
This integral geometric formula expresses the mixed area measure $S_{1}(K,\mathbb{D},{}\cdot{})$ as an average of surface area measures of two-dimensional projections of $K$. The core idea of our argument is to compare ˜2.1 to the disintegration ˜1.5 of the measure $\mu$ to relate the anisotropic Christoffel problem with the disk to the classical Christoffel problem in two dimensions.
By ˜1.1, we explicitly solve the $2$-dimensional Christoffel problem for the measures $\mu_{E}$, giving bodies $K_{E}\in\mathcal{K}(E)$, $E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$. If there is a body $K\in\mathcal{K}(\mathbb{R}^{n})$ such that every $K_{E}$ arises as the orthogonal projection $K|E$ of $K$ onto $E$, then we can conclude $\mu=S_{1}(K,\mathbb{D};{{}\cdot{}})$. However, as the subspaces $E$ intersect only in the line spanned by $e_{n}$, compatibility of the $K_{E}$ is no issue; convexity is asserted by condition ˜(iii) in Theorem˜A.
Let us point here that a similar strategy is promising for the intermediate Christoffel–Minkowski problems with the disk as reference body, replacing ˜2.1 by the $i$-homogeneous counterpart from [Hug2024]. However, in this case, the subspaces do intersect, which directly implies convexity but makes compatibility an issue.
In the following, we first recall some classical facts from convex geometry, in particular on mixed volumes and mixed area measures that will be needed later on. Then we show the existence of the disintegration ˜1.5. In the last section we finally prove Theorem˜A.
2.1. Preliminaries
As a general reference on this section, we refer to the monographs by Gardner [Gardner2006] and Schneider [Schneider2014].
First, recall that the mixed volume $V(K_{1},\dots,K_{n})$, $K_{1},\dots,K_{n}\in\mathcal{K}(\mathbb{R}^{n})$ is defined as the suitable coefficient of the homogeneous polynomial obtained by Minkowski addition
$V(\lambda_{1}K_{1}+\dots+\lambda_{n}K_{n})=\sum_{j_{1},\dots,j_{n}=1}^{n}\lambda_{j_{1}}\cdots\lambda_{j_{n}}V(K_{1},\dots,K_{n}),\qquad\lambda_{1},\dots,\lambda_{n}\geq 0.$
The mixed volume can be represented (or localized) using the Riesz representation theorem,
(2.2) $V(L,K_{1},\ldots,K_{n-1})=\frac{1}{n}\int_{\mathbb{S}^{n-1}}h_{L}(u)\,S(K_{1},\ldots,K_{n-1};du),\quad L\in\mathcal{K}(\mathbb{R}^{n}),$
defining a centered, non-negative Borel measure $S(K_{1},\ldots,K_{n-1};{{}\cdot{}})$ on $\mathbb{S}^{n-1}$, called the mixed area measure of $K_{1},\dots,K_{n-1}\in\mathcal{K}(\mathbb{R}^{n})$. Moreover, by definition, it is $1$-homogeneous and translation-invariant in every component, as well as symmetric under permuting the entries. For $K_{1}=\dots=K_{i}=K$ and $K_{i+1}=\dots=K_{n-1}=B^{n}$, $0\leq i\leq n$, and re-normalizing, we obtain the $i$th intrinsic volume of $K\in\mathcal{K}(\mathbb{R}^{n})$,
(2.3) $V_{i}(K)=\frac{\binom{n}{i}}{\kappa_{n-i}}V(K^{[i]},(B^{n})^{[n-i]}).$
Here, we denote by $K^{[i]}$ the $i$-tuple $(K,\dots,K)$ with $K$ repeated $i$-times.
If all the bodies $K_{1}\ldots,K_{n-1}$ are contained in the hyperplane $u^{\perp}$, then
(2.4) $S(K_{1},\ldots,K_{n-1};{}\cdot{})=V^{u^{\perp}}\!(K_{1},\ldots,K_{n-1})(\delta_{u}+\delta_{-u}).$
Next, recall that mixed area measures are locally determined. Indeed, $S(K_{1},\dots,K_{n-1};\beta)$ depends only on $\tau(K_{1},\beta),\dots,\tau(K_{n-1},\beta)$, where the reverse spherical image $\tau(K,\beta)$ of $K\in\mathbb{R}^{n}$ at a Borel set $\beta\subseteq\mathbb{S}^{n-1}$ is defined as
$\tau(K,\beta)=\bigcup_{u\in\beta}F(K,u),$
with $F(K,u)=\{x\in\mathbb{R}^{n}:h_{K}(u)=\langle x,u\rangle\}$ the face of $K$ in direction $u\in\mathbb{S}^{n-1}$.
Lemma 2.1 ([Schneider2014]*p. 215).
Let $K_{1},K_{1}^{\prime},\ldots,K_{n-1},K_{n-1}^{\prime}\in\mathcal{K}(\mathbb{R}^{n})$ and $\beta\subseteq\mathbb{S}^{n-1}$ be a Borel set such that $\tau(K_{j},\beta)=\tau(K_{j}^{\prime},\beta)$ for all $j\in\{1,\ldots,n-1\}$. Then
$S(K_{1},\ldots,K_{n-1};\beta)=S(K_{1}^{\prime},\ldots,K_{n-1}^{\prime};\beta).$
In order to deal with the uniqueness of the solution $K$ to the Christoffel problem involving the disk, we recall Minkowski’s quadratic inequality, a classical result from Brunn–Minkowski theory. It states that for all convex bodies $K,L,C\in\mathcal{K}(\mathbb{R}^{n})$,
(2.5) $V(K,L,C^{[n-2]})^{2}\geq V(K,K,C^{[n-2]})V(L,L,C^{[n-2]}).$
This inequality was first established by Minkowski [Minkowski1903] in three dimensions and later extended by Bonnesen and Fenchel [Bonnesen1987] to higher dimensions. Its equality conditions were settled only recently in a landmark paper by Shenfeld and van Handel [Shenfeld2022].
For our purposes, we only require the special case where $C$ is lower-dimensional (see [Shenfeld2022]*Theorem 2.3). Here, a vector $u\in\mathbb{R}^{n-1}\setminus\{0\}$ is called a $1$-extremal normal direction of $C$ if there do not exist linearly independent normal vectors $u_{1},u_{2},u_{3}\in\mathbb{R}^{n-1}\setminus\{0\}$ at a boundary point of $C$ such that $u=u_{1}+u_{2}+u_{3}$.
Theorem 2.2 ([Shenfeld2022]).
Let $K,L,C\in\mathcal{K}(\mathbb{R}^{n})$ be such that $C-C\subseteq e_{n}^{\perp}$ and $V(K,L,C^{[n-2]})>0$. Then equality holds in ˜2.5 if and only if $\tilde{L}:=\frac{V(K,L,C^{[n-2]})}{V(L,L,C^{[n-2]})}L$ satisfies that $K+F(\tilde{L},e_{n})$ and $\tilde{L}+F(K,e_{n})$ have the same supporting hyperplanes in all $1$-extremal normal directions of $C$.
Observe now that at every point $x\in\mathbb{D}$, the corresponding normal cone $N(\mathbb{D},x)$ is either one- or two-dimensional. This implies that every direction $u\in\mathbb{S}^{n-1}$ is a $1$-extremal normal direction of $\mathbb{D}$, so ˜2.2 specializes for $C=\mathbb{D}$ to the following statement.
Corollary 2.3.
Let $K,L\in\mathcal{K}(\mathbb{R}^{n})$ be such that $V(K,L,\mathbb{D}^{[n-2]})>0$. Then
$V(K,L,\mathbb{D}^{[n-2]})^{2}=V(K,K,\mathbb{D}^{[n-2]})V(L,L,\mathbb{D}^{[n-2]})$
if and only if $K+F(\tilde{L},e_{n})=\tilde{L}+F(K,e_{n})$, where $\tilde{L}:=\frac{V(K,L,\mathbb{D}^{[n-2]})}{V(L,L,\mathbb{D}^{[n-2]})}L$.
2.2. Disintegration of measures
Next, we want to show that under the given assumptions on the pushforward of the measure $\mu$ in Theorem˜A, there exists a disintegration ˜1.5. To this end, we recall the disintegration theorem from measure theory (see, e.g., [Ambrosio2008]*Theorem 5.3.1).
Theorem 2.4 ([Ambrosio2008]).
Let $X$, $Y$ be two Polish spaces, $\pi:X\to Y$ be a Borel map, $\mu$ be a non-negative Borel measure on $X$, and denote $\nu=\pi_{\ast}\mu$. Then there exists a $\nu$-a.e. uniquely determined family $(\mu_{y})_{y\in Y}$ of probability measures on $X$ such that
  1. (i)
    for every Borel set $\beta\subseteq X$, the map $y\mapsto\mu_{y}(\beta)$ is Borel,
  2. (ii)
    for $\nu$-a.e. $y\in Y$, the probability measure $\mu_{y}$ is concentrated on $\pi^{-1}(y)$, and
  3. (iii)
    for every bounded Borel function $f:X\to\mathbb{R}$,
    $\int_{X}f(x)\,\mu(dx)=\int_{Y}\int_{\pi^{-1}(y)}f(x)\,\mu_{y}(dx)\,\nu(dy).$
This family $(\mu_{y})_{y\in Y}$ is called disintegration of $\mu$. We apply this theorem in the instance where $X=\mathbb{S}^{n-1}\setminus\{\pm e_{n}\}$, $Y=\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$, and $\pi:\mathbb{S}^{n-1}\setminus\{\pm e_{n}\}\to\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n}):u\mapsto e_{n}\vee u$ assigns to the point $u$ the unique $2$-plane containing both $e_{n}$ and $u$. Moreover, we endow the Grassmann manifold $\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$ with the unique $\operatorname{O}(n-1)$ invariant probability measure.
Corollary 2.5.
Let $\mu$ be a non-negative Borel measure on $\mathbb{S}^{n-1}\setminus\{\pm e_{n}\}$ and suppose that $\pi_{\ast}\mu$ is absolutely continuous with a continuous density. Then there exists an a.e. uniquely defined family $(\mu_{E})_{E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})}$ of non-negative Borel measures on $\mathbb{S}^{n-1}$ such that
  1. (i)
    for every Borel set $\beta\subseteq\mathbb{S}^{n-1}$, the map $E\mapsto\mu_{E}(\beta)$ is Borel,
  2. (ii)
    for a.e. $E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$, the measure $\mu_{E}$ is concentrated on $\mathbb{S}^{1}(E)\setminus\{\pm e_{n}\}$, and
  3. (iii)
    for every bounded Borel function $f:\mathbb{S}^{n-1}\to\mathbb{R}$,
    (2.6) $\int_{\mathbb{S}^{n-1}\setminus\{\pm e_{n}\}}f(u)\,\mu(du)=\int_{\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})}\int_{\mathbb{S}^{1}(E)\setminus\{\pm e_{n}\}}f(v)\,\mu_{E}(dv)\,dE.$
Proof.
First, we let $\nu=\pi_{\ast}\mu$ and denote by $(\tilde{\mu}_{E})_{E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})}$ the disintegration of $\mu$ according to ˜2.4. By our assumption on $\mu$, we have $\nu(dE)=\rho(E)dE$ with some density $\rho\in C(\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n}))$. Due to the continuity of $\rho$, we may now define $\mu_{E}:=\rho(E)\tilde{\mu}_{E}$ for $E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$. It is then easy to see that the family $(\mu_{E})$ fulfills the assertion. ∎
2.3. Proof of Theorem˜A
Now we are ready to prove the main result of this article, Theorem˜A. In the proof below, we readily identify $\mu$ with its restriction to $\mathbb{S}^{n-1}\setminus\{\pm e_{n}\}$, and denote by $(\mu_{E})$ its disintegration according to ˜2.5.
Proof of Theorem˜A.
To show that the conditions ˜(i) to ˜(iii) are necessary, suppose that there exists a convex body $K\in\mathcal{K}(\mathbb{R}^{n})$ such that $\mu=S_{1}(K,\mathbb{D},{}\cdot{})$. We identify $\mu$ with its restriction to $\mathbb{S}^{n-1}\setminus\{\pm e_{n}\}$, take some function $\tilde{f}\in C(\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n}))$ and let $f:=\tilde{f}\circ\pi$, that is, $f(u)=\tilde{f}(e_{n}\vee u)$. Then the Cauchy–Kubota type formula ˜2.1, combined with ˜2.2, gives
$\int_{\mathbb{S}^{n-1}}f(u)\,S_{1}(K,\mathbb{D},du)=\frac{\kappa_{n-1}}{2}\int_{\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})}\tilde{f}(E)S^{E}_{1}(K|E,\mathbb{S}^{1}(E))\,dE$
$\qquad=\kappa_{n-1}\int_{\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})}\tilde{f}(E)V_{1}(K|E)\,dE,$
and thus, $\pi_{\ast}\mu=\kappa_{n-1}V_{1}(K|E)dE$. Since the expression $V_{1}(K|E)$ is continuous in $E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$, the push-forward measure $\pi_{\ast}\mu$ is absolutely continuous with a continuous density, showing ˜(i). Moreover, due to ˜2.1 and the uniqueness in ˜2.5, its disintegration $(\mu_{E})_{E}$ is given by $\mu_{E}=\frac{\kappa_{n-1}}{2}S_{1}^{E}(K|E,{}\cdot{})$. Hence, for a.e. $E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$, the measure $\mu_{E}$ (as a measure on $\mathbb{S}^{1}(E)$) is centered, showing ˜(ii). To verify the final condition ˜(iii), we let $h$ be the support function of $K$ and combine ˜1.3 and ˜1.1 in dimension $2$ to obtain
$h(u)-\frac{1}{\pi}\int_{\mathbb{S}^{1}(e_{n}\vee u)}h(v)\langle u,v\rangle\,dv=h_{K|(e_{n}\vee u)-s(K|(e_{n}\vee u))}(u)$
$\qquad=\frac{2}{\pi\kappa_{n-1}}\int_{\mathbb{S}^{1}(e_{n}\vee u)}\sqrt{1-\langle u,v\rangle^{2}}(\pi-\arccos\langle u,v\rangle)\,\mu_{e_{n}\vee u}(dv).$
Conversely, suppose that $\mu$ is a non-negative finite Borel measure on $\mathbb{S}^{n-1}$ satisfying the conditions ˜(i) to ˜(iii), with $(\mu_{E})_{E}$ denoting its disintegration according to ˜2.5. Moreover, take $K\in\mathcal{K}(\mathbb{R}^{n})$ to be the convex body that has $h$ as its support function. For a.e. subspace $E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$, the measure $\mu_{E}$ is centered, so by ˜1.1, there exists a convex body $K_{E}\in\mathcal{K}(E)$ such that $\mu_{E}=S_{1}^{E}(K_{E},{}\cdot{})$. Combining ˜1.3 and ˜1.2, we obtain that
$h_{K|E-s(K|E)}(u)=\int_{\mathbb{S}^{1}(E)}\sqrt{1-\langle u,v\rangle^{2}}(\pi-\arccos\langle u,v\rangle)\,\mu_{E}(dv)=\pi h_{K_{E}-s(K_{E})}(u)$
for every $u\in\mathbb{S}^{1}(E)$. This shows that $K|E-s(K|E)=\pi(K_{E}-s(K_{E}))$ and thus, $\mu_{E}=S_{1}^{E}(K_{E},{}\cdot{})=\frac{1}{\pi}S_{1}^{E}(K|E,{}\cdot{})$. As $\mu_{E}$ has no mass at $\pm e_{n}$, so has $S_{i}^{E}(K|E,{{}\cdot{}})$. Therefore, combining ˜2.1 and ˜2.6, we have that for every bounded Borel function $f:\mathbb{S}^{n-1}\to\mathbb{R}$,
$\int_{\mathbb{S}^{n-1}}f(u)\,\mu(du)=\int_{\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})}\int_{\mathbb{S}^{1}(E)\setminus\{\pm e_{n}\}}f(v)\,\mu_{E}(dv)\,dE$
$\qquad=\frac{1}{\pi}\int_{\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})}\!\int_{\mathbb{S}^{1}(E)\setminus\{\pm e_{n}\}}\!f(u)\,S^{E}_{1}(K|E,du)\,dE=\frac{2}{\kappa_{n-1}\pi}\int_{\mathbb{S}^{n-1}}f(u)\,S_{1}(K,\mathbb{D},du).$
Hence $\mu=\frac{2}{\kappa_{n-1}\pi}S_{1}(K,\mathbb{D},{}\cdot{})$ and rescaling $K$ yields the claim.
For the uniqueness part of the statement, take two convex bodies $K,L\in\mathcal{K}(\mathbb{R}^{n})$ such that $S_{1}(K,\mathbb{D},{}\cdot{})=S_{1}(L,\mathbb{D};{}\cdot{})$, and also, $S_{1}(K,\mathbb{D},\{\pm e_{n}\})=0$. By integrating the respective support function of $K$ and $L$ against these mixed area measures, by ˜2.2, we have that $V(K,K,\mathbb{D}^{[n-2]})=V(K,L,\mathbb{D}^{[n-2]})=V(L,L,\mathbb{D}^{[n-2]})$.
Thus, equality is attained in Minkowski’s quadratic inequality ˜2.5 for $C=\mathbb{D}$, so according to ˜2.3, the bodies $K+F(L,e_{n})$ and $L+F(K,e_{n})$ are equal. Noting, however, that due to Lemma˜2.1 and identities ˜2.3 and ˜2.4,
$S_{1}(K,\mathbb{D};\{\pm e_{n}\})=S_{1}(F(K,\pm e_{n}),\mathbb{D};\{\pm e_{n}\})=\frac{1}{n}V_{1}(F(K,\pm e_{n})),$
the condition $\mu(\pm e_{n})=0$ is equivalent to the faces $F(K,e_{n})$ and $F(K,-e_{n})$ being only singletons. Hence, we conclude that $K$ and $L$ are translates of each other. ∎
Under the addition that $\mu$ is even, which corresponds to $K$ being centrally symmetric, the classification simplifies considerably. This is the content of the theorem below, which we obtain from an easy modification of the proof of Theorem˜A.
Theorem 2.6.
Let $\mu$ be a non-negative even finite Borel measure on $\mathbb{S}^{n-1}$ and $\mu(\{\pm e_{n}\})=0$. Then there exists a convex body $K\in\mathcal{K}(\mathbb{R}^{n})$ such that $\mu=S_{1}(K,\mathbb{D},{}\cdot{})$ if and only if $\pi_{\ast}\mu$ is absolutely continuous with a continuous density and
(2.7) $h(u)=\int_{\mathbb{S}^{1}(e_{n}\vee u)}\sqrt{1-\langle u,v\rangle^{2}}\,\mu_{e_{n}\vee u}(dv),$
defined for a.e. $u\in\mathbb{S}^{n-1}$, is a support function.
Moreover, $K$ is unique up to a translation.
Proof.
Suppose first that there exists a convex body $K\in\mathcal{K}(\mathbb{R}^{n})$ such that $\mu=S_{1}(K,\mathbb{D},{}\cdot{})$. By the evenness of $\mu$ and compatibility of mixed area measures with linear isometries, $\mu=S_{1}(-K,\mathbb{D},{}\cdot{})$. By the uniqueness statement in Theorem˜A, there exists some $x\in\mathbb{R}^{n}$ such that K=xKK=x-K, that is, $K$ is symmetric about $x$.
Due to Theorem˜A, the push-forward measure $\pi_{\ast}\mu$ is absolutely continuous with a continuous density, and from the proof of Theorem˜A, the disintegration of $\mu$ is necessarily given by $\mu_{E}=\frac{\kappa_{n-1}}{2}S_{1}^{E}(K|E,{}\cdot{})$ for a.e. $E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$. In particular, $\mu_{E}$ (as a measure on $\mathbb{S}^{1}(E)$) is even for a.e. $E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$. Moreover, from the proof of Theorem˜A, the support function $h$ of a dilated copy of $K$ satisfies identity ˜1.6. Due to the evenness of $h$ and $\mu_{E}$, this identity then simplifies to ˜2.7.
Conversely, suppose now that $\mu$ is a non-negative finite Borel measure on $\mathbb{S}^{n-1}$ satisfying the conditions stated in the theorem, with $(\mu_{E})_{E}$ denoting its disintegration according to ˜2.5. The evenness of $\mu$ and a.e. uniqueness of its disintegration assert that $\mu_{E}$ (as a measure on $\mathbb{S}^{1}(E)$) is for a.e. $E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$ even, and thus, centered.
Consequently, the right hand side of ˜2.7 is an even function of $u$, and thus, so is the left hand side. Since identity ˜2.7 is simply the even component of identity ˜1.6, we deduce that $h$ also satisfies ˜1.6. In conclusion, $\mu$ meets the requirements of Theorem˜A, and thus, there exists a convex body $K\in\mathcal{K}(\mathbb{R}^{n})$ such that $\mu=S_{1}(K,\mathbb{D},{}\cdot{})$. ∎
We turn to the special case where $\mu$ is absolutely continuous with respect to the spherical Lebesgue measure, that is $\mu(du)=q(u)du$ for some $q\in L^{1}(\mathbb{S}^{n-1})$. We have the following decomposition of the spherical Lebesgue measure into spherical cylinder coordinates at our disposal (see, e.g., [Mueller1966]*p. 1)
$\int_{\mathbb{S}^{n-1}}q(u)\,du=\frac{1}{2}\int_{\mathbb{S}^{n-2}(e_{n}^{\perp})}\int_{\mathbb{S}^{1}(e_{n}\vee u)}q(v)\lvert\langle u,v\rangle\rvert^{n-2}\,dv\,du.$
From this formula, it is immediate that $\pi_{\ast}\mu$ is absolutely continuous with a continuous density given by $\frac{d\pi_{\ast}\mu}{dE}=\frac{\omega_{n-1}}{2}\int_{\mathbb{S}^{1}(E)}q(v)\lVert v|E\rVert^{n-2}\,dv$. Moreover, the disintegration of $\mu$ in the sense of ˜2.6 is given by
$\mu_{E}(dv)=\frac{\omega_{n-1}}{2}q(v)\lVert v|E\rVert^{n-2}\,dv.$
Hence, as a special case of Theorem˜A, we obtain the following.
Corollary 2.7.
Let $q$ be a non-negative $L^{1}(\mathbb{S}^{n-1})$ function. Then there exists a convex body $K\in\mathcal{K}(\mathbb{R}^{n})$ such that $q(u)du=S_{1}(K,\mathbb{D},du)$ if and only if $\int_{\mathbb{S}^{1}(e_{n}\vee u)}q(u)\lvert\langle u,v\rangle\rvert^{n-2}v\,dv=0$ for almost every $u\in\mathbb{S}^{n-2}(e_{n}^{\perp})$, and there exists a support function $h$ such that
$h(u)-\frac{1}{\pi}\int_{\mathbb{S}^{1}(e_{n}\vee u)}h(v)\langle u,v\rangle\,dv=\int_{\mathbb{S}^{1}(e_{n}\vee u)}\sqrt{1-\langle u,v\rangle^{2}}(\pi-\arccos\langle u,v\rangle)\lvert\langle u,v\rangle\rvert^{n-2}q(v)\,dv$
for a.e. $u\in\mathbb{S}^{n-1}$.
Moreover, $K$ is unique up to a translation.
Under the additional assumption of evenness, ˜2.6 yields the following.
Corollary 2.8.
Let $q$ be a non-negative even $L^{1}(\mathbb{S}^{n-1})$ function. Then there exists a convex body $K\in\mathcal{K}(\mathbb{R}^{n})$ such that $q(u)du=S_{1}(K,\mathbb{D},du)$ if and only if
$h(u)=\int_{\mathbb{S}^{1}(e_{n}\vee u)}\sqrt{1-\langle u,v\rangle^{2}}\lvert\langle u,v\rangle\rvert^{n-2}q(v)\,dv,$
defined for a.e. $u\in\mathbb{S}^{n-1}$, is a support function.
Moreover, $K$ is unique up to a translation.
Acknowledgments
The first-named author was supported by the Deutsche Forschungsgesellschaft (DFG), project 520350299. The second-named author was supported by the Austrian Science Fund (FWF), project doi:10.55776/P34446. The third-named author was supported by the Austrian Science Fund (FWF), project doi:10.55776/ESP236, ESPRIT program.